On January 1, Abel was a 25% partner in Corn General Partnership, which had partnership liabilities of $300,000. On January 2, a new partner was admitted and Abel’s interest was reduced to 20%. On April 1, Corn repaid a $100,000 general partnership loan. Ignoring any income, loss, or distributions for the year, what was the net effect of the transactions described for Abel’s tax basis in Corn partnership interest?

Group of answer choices

Has no effect.

Decrease of $35,000.

Increase of $15,000

Decrease of $75,000.
Answer #1

Decrease of $35,000.

Explanation:

Before appointment of Corn

Abel Liability': 300000*25%= 75000

After appointment 200000*20% = 40000

Therefore decrease= 75000-40000= 35000